On the spectrum of curve singularities

by András Némethi

The Ohio State University Columbus, OH 43210, USA.

Dedicated to Egbert Brieskorn on the occasion of his 60th birthday

§1. Introduction

There is the following crucial problem in the singularity theory: how can we recognize that a singular germ is hypersurface, or complete intersection singularity. Or: is there any criterion which distinguishes the isolated hypersurface singularities among the isolated singularities? These kind of questions arise in many areas of the algebraic geometry and singularity theory, we mention here only one.

A. Durfee [3] conjectured that the Milnor fiber of an isolated complete intersection singularity of dimension two is negative. Although J. Wahl [23] found a smoothing of a (non-hypersurface) singularity with positive signature, there is still a strong belief that the conjecture is true for hypersurfaces. But, if the conjecture is true, then what is special in hypersurfaces, which makes their signature negative?

In this paper we try to understand this question for hypersurfaces of type f(x, y) + z^{N}. In [12, 10], we give a series of properties of the resolution graphs of germs f : (X, x) → (C, 0), defined on a normal surface singularity (X, x), which distinguish the plane curve singularities. Here, we present some properties ({P(N)}_{2≤N≤∞}) of the spectrum, which codifies a part of the information of the mixed Hodge structure on the vanishing cohomology of the germs. The property P(N), for plane curve singularities is equivalent to the negativity of the equivariant signature σ_{≠1}(f + z^{N}) of the germ f(x, y) + z^{N}. We prove that these properties are valid for plane curve singularities (in particular, we prove Durfee’s conjecture for hypersurfaces of type f(x, y) + z^{N}) (cf. 5.2), and we give germs f : (X, x) → (C, 0) (with (X, x) non-smooth) which do not satisfy any of the properties P(N).

In the paper we emphasize the arithmetical flavour of the spectrum and its connections with number theory, namely with generalized Dedekind sums.

§2. The properties P(N) (2 ≤ N ≤ ∞).

We start this section with some definitions (their motivation will come later, cf. 2.7). 2.1. Definitions. Let N^{Q} be the free abelian monoid generated by Q: its elements are finite sums of the form ∑_{r∈Q} n_{r}(r), where n_{r} ∈ N = {0, 1, …}. We will use the notation S_{N} for ∑_{1≤k<N}(k/N) (N ≥ 2). If S = ∑_{r∈(0,1]} n_{r}(r), we define the symmetric element with respect r = 1 by S̃ = n_{1}(1)+∑_{r∈(0,1)}(n_{r}(r)

+n_{r}(2 − r)). The degree of S = ∑ n_{r}(r) is given by μ(S) = ∑ n_{r} ∈ N, the “signature of S” by σ(S) = ∑_{r} n_{r}sign sin πr = ∑_{r∉Z}(−1)^{[r]}n_{r} (where [·] denotes the integer part), and

m(S) = ∑_{r} n_{r} ∑_{k∈Z}(−1)^{k}m([r, 1 + r] ∩ [k, k + 1]),

where, for an interval I ⊂ R, m(I) denotes its length.

The multiplicative structure of N^{Q} is generated by (r) ⊗ (s) = (r + s). Then it is easy to see that

m(S) = lim_{N→∞} \frac{σ(S ⊗ S_{N})}{N}.

2.2. Definition.

a) Fix an element S = ∑_{r∈(0,1]} n_{r}(r) and an integer N ≥ 2. We say that S satisfies the property P(N) if σ(S̃ ⊗ S_{N}) < 0.

b) An element S = ∑_{r∈(0,1]} n_{r}(r) satisfies the property P(∞) if m(S̃) < 0. 2.3. Remark. If S = ∑_{r∈(0,1)} n_{r}(r), then σ(S̃ ⊗ S_{N}) = 2σ(S ⊗ S_{N}) and similarly m(S̃) = 2m(S). Therefore, P(N) (resp. P(∞)) is equivalent to σ(S ⊗ S_{N}) < 0 (resp. m(S) < 0).

We are really interested in elements S provided by the spectrum of an isolated curve singularity. 2.4. Example. Let (X, x) be a normal surface singularity, with the property that its link K_{X} is rational homology sphere (i.e. H_{1}(K_{X}, Q) = 0). (This condition is not really necessary, but it will simplify our presentation. For the general situation, see for example, [14].) Let f : (X, x) → (C, 0) be a germ of an analytic function which defines a one–dimensional isolated singularity. We consider an embedded resolution ϕ : (Y, D) → (X, f^{−1}(0)) of (f^{−1}(0), x) ⊂ (X, x) (here D = ϕ^{−1}(f^{−1}(0))). Let E = ϕ^{−1}(x) be the exceptional divisor and E = ∪_{w∈W}E_{w} be its decomposition in irreducible divisors. All these irreducible divisors are rational. Let ∪_{a∈A}S_{a} be the irreducible decomposition of the strict trasform S of f^{−1}(0). Then D = E ∪S. Let G_{f} be the resolution graph of f, i.e. its vertices V = W ∐ A consist of the nonarrowhead vertices W (corresponding to the irreducible exceptional divisors), and arrowhead vertices A (correponding to the strict transform divisors of D). We will assume that no irreducible exceptional divisor has an self intersection and W ≠ ∅. If two irreducible divisors corresponding to v_{1}, v_{2} ∈ V have an intersection point then (v_{1}, v_{2}) (= (v_{2}, v_{1})) is an edge of G_{f}. The set of edges is denoted by E.

For any w ∈ W, we denote by V_{w} the set of vertices v ∈ V adjacent to w. The graph G_{f} is decorated by the self intersection (or Euler-) numbers

e_{w} := E_{w} · E_{w} for any w ∈ W.

For any v ∈ V let m_{v} be the multiplicity of f ∘ ϕ along the irreducible divisor corresponding to v. In particular, for any a ∈ A one has m_{a} = 1. The multiplicities satisfy the following relations. For any w ∈ W one has:

(2.5) e_{w}m_{w} + ∑_{v∈V_{w}} m_{v} = 0.

These relations determine the multiplicities {m_{w}}_{w∈W} in terms of the autointersection numbers {e_{w}}_{w}. For an edge e = (v_{1}, v_{2}) define m_{e} = g.c.d.(m_{v_{1}}, m_{v_{2}}).

The vanishing cohomology H^{1}(F, C) of the Milnor fiber F of f has a natural mixed Hodge structure. The associated spectrum Sp(f) is in the interval (−1, 1), and it is symmetric with respect to r = 0 (for details, see for example, [21, 20, 14]). We define S(f) = ∑(r + 1), where the sum is over the spectral elements r of Sp(f) which are in the interval (−1, 0]. Obviously S̃(f) is exactly Sp(f) shifted one unit to the right. In particular μ(S̃(f)) = dim H^{1}(F, C).

The element S(f) can be computed from the graph G_{f} as follows; 2.6. Proposition. [20, 14] Define R_{w}^{s} := ∑_{v∈V_{w}}{sm_{v}/m_{w}}, where {·} denotes the fractional part. Then:

S(f) = (#A − 1)(1) + ∑_{e∈E} ∑_{0<s<m_{e}}(\frac{s}{m_{e}}) + ∑_{w∈W} ∑_{0<s<m_{w}}(−1 + R_{w}^{s})(\frac{s}{m_{w}}).

2.7. Remark. The relation between the above example and the properties P(N) is the following. If g : (C^{3}, 0) → (C, 0) is an isolated singularity, then its equivariant signature σ_{λ} for λ ≠ 1 can be computed from its spectrum Sp(g) only (for the exact formula, see e.g. [9, 11]). In general σ_{1} cannot be computed from Sp(g) alone (cf. 6.10. [11]). If g(x, y, z) = f(x, y) + z^{N}, then Sp(g), shifted by one unit, is exactly S̃(f) ⊗ S_{N} (cf. [18, 22]); and σ_{≠1}(f + z^{N}) = σ(S̃(f) ⊗ S_{N}). The equivariant signature σ_{1}(f + z^{N}) is zero if the monodromy h_{1}(g) corresponding to the eigenvalue one has no Jordan blocks of size two. In particular, (cf. [6]) if f is irreducible, then σ_{1}(g) = 0. In general σ_{1}(g) = h_{1}^{22}(g) ≥ 0 (cf. 6.7. [11]), hence σ(S̃(f)⊗S_{N}) ≤ σ(f +z^{N}).

Now, the conjecture of Durfee [3], mentioned in the introduction, says that the signature of any hypersurface singularity g : (C^{3}, 0) → (C, 0) is negative. Property P(N) (2 ≤ N < ∞) for S(f) corresponds exactly to this conjecture. The function N ↦ σ(S̃(f) ⊗ S_{N}) is a sum of a periodic and a linear function (say P (N) and c · N); (cf. 4.2). Hence P(∞) is equivalent to c < 0.

We expect that the above results (valid in the hypersurface situation) have the corresponding analogues in the general case, (i.e. when f is not plane curve singularity). On the other hand, the properties P(N) make sense for arbitrary germ f : (X, 0) → (C, 0). The goal of the present paper is to show that the properties P(N) distinguish the plane curve singularities among the general germs f : (X, 0) → (C, 0).

We end this section with the following remark: for any fixed 2 ≤ N_{1} < N_{2} < ∞, the properties P(N_{1}), P(N_{2}) and P(∞) are independent. 2.8. Example. Consider a sufficiently small 1 >> r > 0.

a) S = 2(1/2 − r) + (1 − r) satisfies P(∞) and P(3), but not P(2);

b) S = (r) + 2(1/2 + r) satisfies P(2), but not P(3) and P(∞).

§3. Positive results. The case of plane curve singularities.

In this section we study the set S(f), where f : (C^{2}, 0) → (C, 0) is an isolated singularity.

I. Combinatorial–arithmetical approach

Since S(f) is constant under the μ–constant deformations, S(f) depends only on the topological type of the embedding (f^{−1}(0), 0) ⊂ (C^{2}, 0) (this fact can be read also from 2.6). Moreover, in the irreducible case, the topological invariants of this embedding behave rather additively with respect to the splice decomposition of the of f^{−1}(0) (or of the graph G_{f}) (see, for example, [4]). Hence, one can expect that for irreducible f the invariant S(f) have a nice representation in terms of the Newton pairs of f. In the Brieskorn case f(x, y) = x^{p} +y^{q}, ((p, q) = 1) one has S(f) = ∑(l/p+k/q), where the sum is over 0 < l < p, 0 < k < q, l/p+k/q < 1. It is convenient to use the notation S(p, q) for this element of N^{Q}. One can expect that for general irreducible f, S(f) can be obtained by “iterating” the Brieskorn case. The following theorem provides such a formula (cf. 2.2 in [20], and an unpublished paper of M. Saito). 3.1. Theorem. Assume that f : (C^{2}, 0) → (C, 0) is irreducible with Newton pairs (p_{i}, q_{i})_{i=1}^{r}. Define the integers {a_{i}}_{i=1}^{r} by

a_{1} = q_{1} and a_{i+1} = q_{i+1} + p_{i+1}p_{i}a_{i} if i ≥ 1.

(These numbers are the decorations of the splice diagram of f, cf. [4].) Then:

S(f) = ∑_{i=1}^{r} S_{i}, where S_{i} = ∑(\frac{k/a_{i} + l/p_{i} + t}{p_{i+1}p_{i+2} ⋯ p_{r}}),

where the second sum is over 0 < k < a_{i}, 0 < l < p_{i}, k/a_{i} + l/p_{i} < 1 and 0 ≤ t ≤ p_{i+1} ⋯ p_{r} − 1 (if i = r, then S_{r} = S(a_{r}, p_{r})). Proof. If #V_{w} > 2, then w is called “rupture point”. Consider the relation given by (2.6). If #V_{w} = 1, then (by 2.5) R_{w}^{s} = 0, if #V_{w} = 2, then R_{w}^{s} = 1 iff m_{w}|sm_{v} (v ∈ V_{w}). By an elementary computation one can show that S(f) is concentrated only in the rupture points {w_{1}, …, w_{r}} of G_{f}, namely:

(3.2) S(f) = ∑_{i=1}^{r} ∑_{{0<s<mw_{i} \atop mw_{i} |̸sd_{i}}} (−1 + ∑_{v∈V_{w_{i}}}\{sm_{v}/m_{w_{i}}\})(\frac{s}{m_{w_{i}}}),

where d_{i} := g.c.d.((m_{v})_{v∈V_{w_{i}}}, m_{w_{i}}) (cf. [20] and [19]).

Now, m_{w_{i}} = a_{i}p_{i}p_{i+1} ⋯ p_{r} (for details, see e.g. [4]). Consider the integers b_{i} and c_{i} such that b_{i}p_{i} + c_{i}a_{i} = 1. Then the multiplicities {m_{v}} (v ∈ V_{w_{i}}), modulo m_{w_{i}}, are: (−b_{i}p_{i}p_{i+1} ⋯ p_{r}, −c_{i}a_{i}p_{i+1} ⋯ p_{r}, p_{i+1} ⋯ p_{r}) (cf. [15, 19]). Hence: d_{i} = p_{i+1} ⋯ p_{r}. Therefore R_{w_{i}}^{s} = {−b_{i}s/a_{i}} + {−c_{i}s/p_{i}} + {s/a_{i}p_{i}}. Now, write s = kp_{i} + la_{i} + ta_{i}p_{i}, where 0 ≤ k < a_{i}, 0 ≤ l < p_{i}, 0 ≤ t < p_{i+1} ⋯ p_{r}. Then m_{w_{i}}|sd_{i} iff (k, l) = (0, 0), and R_{w_{i}}^{s} = {−k/a_{i}} + {−l/p_{i}} + {k/a_{i} + l/p_{i}}, hence the result follows. □

By (3.1) the Milnor number μ(f) of f is 2μ(S(f)) = 2 ∑_{i} μ(S_{i}), where 2μ(S_{i}) = (a_{i}−1)(p_{i}−1)p_{i+1} ⋯ p_{r}. The next theorem gives a similar additivity formula for the signature σ(f + z^{N}) of f(x, y) + z^{N}.

Define ((x)) by {x} − 1/2 if x ∉ Z, and ((x)) = 0 otherwise. 3.3. Theorem. Let f be as in (3.1), and D_{i} := g.c.d.(N, p_{i+1} ⋯ p_{r}) for 1 ≤ i < r, and D_{r} = 1. Then

σ(S(f) ⊗ S_{N}) = ∑_{i=1}^{r} D_{i} · σ(S(a_{i}, p_{i}) ⊗ S_{N/D_{i}}).

Proof. First we consider two lemmas. 3.4. Lemma. Let α ∈ (0, 1), and N ≥ 2 an integer. Then:

σ((α) ⊗ S_{N}) = 2((Nα)) − 2N((α)).

The proof is easy and it is left to the reader (write i/N ≤ α < (i + 1)/N). 3.5. Lemma.

∑_{t=0}^{P −1}((N · \frac{x + t}{P})) = (N, P) · ((\frac{N}{(N, P)} · x)).

Proof. Use the Fourier sine expansion:

((x)) = −\frac{1}{π} ∑_{k=1}^{∞} \frac{1}{k} sin 2πkx,

(or an elementary, longer computation). □ Now (3.4) and (3.5) applied for x = k/a_{i} + l/p_{i} gives: σ(S_{i} ⊗ S_{N}) = D_{i} · σ(S(a_{i}, p_{i}) ⊗ S_{N/D_{i}}), which ends the proof. □ The discussion (2.7) gives: 3.6. Corollary. Let f be as in (3.1), and D_{i} as in (3.3). Then:

σ(f + z^{N}) = ∑_{i=1}^{r} D_{i} · σ(x^{a_{i}} + y^{p_{i}} + z^{N/D_{i}}).

Now, for the Brieskorn singularities it is well known that σ(x^{a} + y^{p} + z^{N}) < 0, and actually σ(x^{a} + y^{p} + z^{N}) < (a − 1)(p − 1)(N − 1)/3, therefore lim_{N→∞} σ(x^{a} + y^{p} + z^{N})/N < 0 as well (cf. 4.3). This gives:

3.7. Corollary. For any irreducible plane curve singularity f, the element S(f) satisfies the properties P(N), (2 ≤ N ≤ ∞).

Unfortunately, the above arithmetical results (i.e. 3.1 and 3.6) have no analogs in the reducible case (at least known by the author). In the next subsection, we will prove the analog of (3.7) for reducible plane curves.

II. Geometrical approach

We start with the following theorem, which supports the belief that Durfee’s conjecture ([3]) is true for hypersurface singularities. 3.8. Theorem. Assume that the germ f : (C^{2}, 0) → (C, 0) defines an isolated singularity and it has r irreducible components. Then the signature of g(x, y, z) = f(x, y) + z^{N} satisfies:

σ(f + z^{N}) ≤ −(N − 1)(r − 1) ≤ 0.

Proof. We start with the following remark. Assume that g : (C^{3}, 0) → (C, 0) defines an isolated singularity with Milnor lattice L(g). If g_{t} is a deformation of g such that g_{0} = g and g_{t}, for t ≠ 0 small, has k singular points with Milnor lattices L_{1}, …, L_{k}, then there is a lattice embedding ⊕_{i=1}^{k}L_{i} ↪ L(g). If c is the codimension of the embedding, then: (∗) σ(g) ≤ c+ ∑_{i} σ(L_{i}). Now consider a deformation {f_{t}}_{t} of f such that f_{0} = f and f_{t}, for t ≠ 0 small, has exactly δ(f) nodes, where μ(f) = 2δ(f)−r+1 (cf. [7]). The deformation g_{t} = f_{t}+z^{N} gives an embedding of δ(f) copies of L(x^{2}+y^{2}+z^{N}) in L(f +z^{N}) with codimension c = (N − 1)(μ − δ), hence (∗) gives the wanted inequality. □ Since σ(S̃(f) ⊗ S_{N}) ≤ σ(f + z^{N}) (cf. 2.7), (3.7) and (3.8) gives: 3.9. Corollary. For any plane curve singularity f, the element S(f) satisfies the properties P(N), (2 ≤ N ≤ ∞).

§4. The arithmetical approach revisited. Dedekind sums.

We will consider again a germ f : (X, x) → (C, 0) as in (2.4). For simplicity, we will assume that the graph G_{f} has only one rupture point w.

Since in (2.6) the contribution of the non–rupture points is cancelled with the contribution of the edges, one has:

(4.1) S(f) = (#A − 1)(1) + ∑_{0<s<m_{w}}(−1 + ∑_{v∈V_{w}} \{sm_{v}/m_{w}\})(\frac{s}{m_{w}}).

Using (3.4) and ∑_{0<s<b}((as/b)) = 0 (apply 3.5 for x = 0), one has:

σ(S̃(f)⊗S_{N}) = −(#A−1)(N−1)+4· ∑_{0<s<m_{w}} ∑_{v∈V_{w}} ((\frac{sm_{v}}{m_{w}}))(((\frac{Ns}{m_{w}}))−N((\frac{s}{m_{w}}))).

We recall that the generalized Dedekind sum (cf. [17, 24]) is defined by:

s(b, c; a) = ∑_{0<s<a}((\frac{sb}{a})) · ((\frac{sc}{a})) (where a > 0).

Notice that s(b, 1; a) = s(1, b; a) is exactly the classical Dedekind sum. 4.2. Corollary. Let (X, x) and f be as above. Then:

σ(S̃(f)⊗S_{N}) = −(#A−1)(N −1)+4· ∑_{v∈V_{w}} (s(m_{v}, N; m_{w})−N ·s(m_{v}, 1; m_{w})),

and:

m(S̃(f)) = −(#A − 1) − 4 ∑_{v∈V_{w}} s(m_{v}, 1; m_{w}).

In particular, the properties P(N) are related to the arithmetical properties of the Dedekind sums (associated with the multiplicity structure of G_{f}). 4.3. Example. Consider again f : (C^{2}, 0) → (C, 0), f(x, y) = x^{p} + y^{a}, (a, p) = 1, a > 1, p > 1. Fix b and c such that bp + ca = 1. Then m_{w} = ap and {m_{v}} = {−bp, −ca, 1} (cf. the proof of 3.1). We recall the following properties of the Dedekind sums:

s(b, c; a) = (a, b, c) · s(\frac{b}{(a, b)}, \frac{c}{(a, b, c)}; \frac{a}{(a, b)}), and :

s(b, c; a) = s(kb, kc; a) if (k, a) = 1.

Using these identities one has: m(S̃(f))/4 = s(p, 1; a)+s(a, 1; p)−s(1, 1; ap). Finally, the famous Dedekind’s Reciprocity Law says that:

s(b, 1; a) + s(a, 1; b) = −\frac{1}{4} + \frac{a^{2} + b^{2} + (a, b)^{2}}{12ab} .

Therefore: m(S̃(f)) = −(a^{2} − 1)(p^{2} − 1)/(3ap) < 0, in particular:

σ(x^{a} + y^{p} + z^{N}) = −N \frac{(a^{2} − 1)(p^{2} − 1)}{3ap} + 4 ∑_{v∈V_{w}} s(m_{v}, N; ap).

Therefore our Corollary (4.2) gives, in this very particular case, (via Brieskorn formula of the signature [2]), the number of lattice points in the tetrahedron (0, 0, 0), (0, 0, a), (0, p, 0), (N, 0, 0) in terms of Dedekind sums. This problem is well–known in number theory, it was solved by Mordell [8] (in the case when a, p, N are relatively prime numbers), and recently by Pommersheim [16] (in the general case). We invite the reader to complete our example for the case (a, p) ≠ 1 (in order to provide a new proof of Pommersheim’s formula, cf. [12]).

§5. Germs without property P(N).

5.1. Fix an integer k ≥ 1. Consider the graph G(k):

[figure: …; …; (1); (1); (1); (2); −k − 1; −2; −2; −2; −2; −2; −2; −2; (2); (3); (2k); (2k + 1); (2k); (3); (2)]

One can verify that the intersection matrix of the graph is negative definite (here it is useful the method given in [4], page 154). Then by a theorem of Grauert [5], there exists a germ f_{k} : (X_{k}, x) → (C, 0) as in (2.4), such that G_{f_{k}} = G(k). Then, (4.1) gives:

S(f) = 2(1) + ∑_{1≤s≤k}(\frac{s}{2k + 1}),

therefore: m(S̃(f_{k})) = −2 + 2 ∑_{1≤s≤k}(1 − \frac{2s}{2k + 1}) = −2 + \frac{2k^{2}}{2k + 1}.

This shows that there is no upper bound for m(S̃(f)), in general. By (4.2):

σ_{N}^{(k)} := σ(S̃(f_{k}) ⊗ S_{N}) = 2 + m(S̃(f_{k})) · N + P^{(k)}(N),

where P^{(k)}(N) is an odd periodic function in N, of period 2k + 1.

If k = 2, then σ_{N}^{(2)} = 2 − 2N/5 + P^{(2)}(N), where for N = 0,; ±1; ±2 one has: P^{(2)}(N) = 0; ∓8/5; ±4/5. In particular, σ_{N}^{(2)} ≥ 0 iff N = 2, 3, 4, 5, 7, 9.

If k = 3, then σ_{N}^{(3)} = 2 + 4N/7 + P^{(3)}(N), where for N = 0; ±1; ±2; ±3 one has: P^{(3)}(N) = 0; ∓18/7; ±6/7; ±2/7. In particular, σ_{N}^{(3)} > 0 for any N ≥ 2.

The case N = 2 also can be computed explicitly. By the Dedekind’s Reciprocity Law one has:

s(1, 1; 2k + 1) = s(2, 2; 2k + 1) = \frac{k(2k − 1)}{6(2k + 1)}, s(2, 1; 2k + 1) = \frac{k(k − 2)}{6(2k + 1)}.

Therefore, (4.2) gives: σ_{2}^{(k)} = −2 + 2k ≥ 0.

Now, assume that k ≥ 4 and N ≥ 3. Then by the inequality:

|s(b, c; a)| ≤ \frac{(a − 1)(a − 2)}{12a}; (a > 0),

(cf. [12, 13]), and by (4.2):

σ_{N}^{(k)} ≥ 2 + N(−2 + \frac{2k^{2}}{2k + 1}) − \frac{2k(2k − 1)}{2k + 1} > 0.

Therefore, for k ≥ 3, S(f_{k}) does not satisfy the property P(N) for any 2 ≤ N ≤ ∞, and for k = 2, P(N) is satisfied only if N ∉ {2, 3, 4, 5, 7, 9}.

5.2. Final Remarks.

a). Actually, for any c ∈ R, we can define similar properties as above: we say that S = ∑_{r∈(0,1]} n_{r}(r) satisfies P(N, c) if σ(S̃⊗S_{N}) < c·μ(S̃⊗S_{N}). The case c = −1/3 is optimal: if c < −1/3, then for sufficiently large d, P(d, c) is not true for S(x^{d} + y^{d}). By another conjecture of Durfee ([3]), P(N, −1/3) is true for any plane curve singularity f.

b). For a plane curve singularity f, the negativity of the signature of f + z^{N} was firstly proved by Ashikaga [1] in a very long paper with a rather sophisticated method. Actually, he proved that the order of σ(f + z^{N}) is less than −N · m(f)^{2}/3, where m(f) is the multiplicity of f. In [12, 13], by a more conceptual approach, we reprove these facts, together with some stronger results. For example, in [12] we prove that P(N, −1/3) is true for any N and any irreducible plane curve singularity f. In [13] we consider the case when f is not irreducible, and we prove that P(N, −1/3) is true with the following restrictions: we assume that if N is smaller than a bound B(f) then λ = 1 is not an eigenvalue of the monodromy of f + z^{N}.

References

[1] Ashikaga, T.: The Signature of the Milnor fiber of Complex Surface Singularities on Cyclic Coverings, preprint, (1995).

[2] Brieskorn, E.: Beispiele zur Differentialtopologie von Singularitäten, Invent. Math., 2., 1-14 (1966).

[3] Durfee, A.: The Signature of Smoothings of Complex Surface Singularities, Math. Ann., 232, 85-98 (1978).

[4] Eisenbud, D. and Neumann, W.: Three-Dimensional Link Theory and Invariants of Plane Curve Singularities, Ann. of Math. Studies 110, Princeton University Press, 1985.

[5] Grauert, H.: Über Modifikationen und exzeptionelle analytische Mengen, Math. Ann., 146, 331–368 (1962).

[6] Lê Dũng Tráng: Sur les noeuds algébriques, Compositio Math., 25, 281-321 (1972).

[7] Milnor, J.: Singular Points of Complex Hypersurfaces, Annals of Math. Studies, Vol. 61, Princeton University Press, 1868.

[8] Mordell, L.J.: Lattice points in a tetrahedron and generalized Dedekind sums, J. Indian Math., 15, 41-46 (1951).

[9] Némethi, A.: The equivariant signature of hypersurface singularities and eta– invariant, Topology, 34, 243-259 (1995).

[10] Némethi, A.: The eta–invariant of variation structures I, Topology and its Applications, 67, 95-111 (1995).

[11] Némethi, A.: The real Seifert form and the spectral pairs of isolated hypersurface singularities, Compositio Math., 98, 23-41 (1995).

[12] Némethi, A.: Dedekind sums and the signature of f(x, y) + z^{N}, submitted.

[13] Némethi, A.: Dedekind sums and the signature of f(x, y)+z^{N}, II, submitted.

[14] Némethi, A. and Steenbrink, J.: Spectral pairs, mixed Hodge modules and series of plane curve singularities, New York Journal of Math., August 16 (1995) (http://nyjm.albany.edu:8000/j/v1/Nemethi-Steenbrink.html)

[15] Neumann, W.: Splicing Algebraic Links, Advanced Studies in Pure Math., 8, 349-361 (1986). (Proc U.S.-Japan Seminar on Singularities 1984)

[16] Pommersheim, J. E.: Toric varieties, lattice points and Dedekind sums, Math. Ann., 295, 1-24 (1993).

[17] Rademacher, H.: Generalization of the Reciprocity formula for Dedekind sums, Duke Math. Journal, 21, 391-397 (1954).

[18] Scherk, J. and Steenbrink, J. H. M.: On the Mixed Hodge Structure on the Cohomology of the Milnor Fiber, Math. Ann., 271, 641-665 (1985).

[19] Schrauwen, R.: Topological Series of Isolated Plane Curve Singularities, Eiseignement Mathématique, 36, 115-141 (1991).

[20] Schrauwen, R., Steenbrink, J. and Stevens, J.: Spectral Pairs and Topology of Curve Singularities, Proc. Sympos. Pure Math., 53, 305-328 (1991).

[21] Steenbrink, J. H. M.: The spectrum of hypersurface singularities, Asterisque, 179-180, 211-223 (1989.

[22] Varchenko, A. N.: Asymptotic Hodge structure in the vanishing cohomology, Math. USSR Izv., 18, 469-512 (1982).

[23] Wahl, J.: Smoothings of normal surface singularities, Topology, 20, 219-246 (1981).

[24] Zagier, D.: Higher dimensional Dedekind sums, Math. Ann., 202, 149-172 (1973).